I need help on solving a linear system using elimination.
The question is
a/2 + b/3 = 1
a/4 - 2b/3 = -1
Thank you very much! (I may need your help for many things in math as well!:))

Answer by Alan3354(69443)

a/2 + b/3 = 1
a/4 - 2b/3 = -1
-------------
Step 1, eliminate fractions.
a/2 + b/3 = 1 times 6 --> 3a + 2b = 6
a/4 - 2b/3 = -1 time 12 --> 3a - 8b = -12
-------------
3a + 2b = 6
3a - 8b = -12
---------------- Subtract
0a + 10b = 18
b = 18/10 = 9/5
------------------
Sub for b in either eqn (with integers) and find a.